find the equation of the inverse of the following functions
f(x)=x^3+5
y=x^3+5
y-5=x^3
x=(y-5)
f(y)=(y-5)
f(x)=(x-5)
can you tell me if this is correct.
Thanks

Answer by Cromlix(4381)

f(x) = x^3 + 5
y = x^3 + 5
Change x for y and y for x
x = y^3 + 5
x - 5 = y^3
y = cube root of(x - 5)
Hope this helps.
:-)